How to refer to a source or designate a new one from the IAU Commission 5 Working Group on Designations

(for astronomical sources of radiation outside the solar system)

(current chair, Marion Schmitz)

(version 25 February 2000)

In general all source listings should always contain positional information and/or a second designation next to a principal designation in order to avoid ambiguities which can arise with a single designation.

Existing designations when used in listings, should never be altered and a bibliographical reference for the designation should be given.

When creating a new catalogue of sources. please consider registering for an acronym, preferably before your new objects become referenced (even informally).

When creating new designations, such as for newly discovered source(s), the following procedure is recommended:

1. Think up an acronym of at least three characters and check the on–line Interactive Dictionary of Acronyms to be sure your acronym is unique.

2. Think up a sequence numbering

(a) running number, or

(b) based on the position, e.g. a chain based on truncated 2000 coordinates preceded by the letter J (example: J1234+5612)

• Other examples and rules are in the IAU Recommendations for Nomenclature, including helpful hints.

• Consult representatives of the Clearing House if problems arise and specific advice is desired.

References

• IAU Recommendations for Nomenclature is revised every three years;

• Interactive Dictionary of Acronyms is updated on a regular basis (i.e. weekly)

The IAU Recommendations for Nomenclature is also available on the World-Wide-Web via your favorite web browser at the URL address: http://cdsweb.u-strasbg.fr/iau-spec.htx

World-Wide-Web access to the Interactive Dictionary of Acronyms is managed by the Centre de Données de Strasbourg (CDS) at the URL address: http://cdsweb.u-strasbg.fr/cgi-bin/Dic; mirror installations are available at http://vizier.nao.ac.jp/vizbin/Dic http://vizier.cfa.harvard.edu/viz-bin/Dic

Latex and PDF versions of How to refer to a source or designate a new one and IAU Recommendations for Nomenclature are accessible by anonymous ftp at Internet node cdsarc.u-strasbg.fr. (URL ftp://cdsarc.u-strasbg.fr/pub/iau)

Clearing House Advice on specific problems may be obtained from members of the Clearing House, a subset of the Working Group on Astronomical Designations of IAU Commission 5 (current chair, Marion Schmitz) Representatives are listed at the end of the IAU Recommendations for Nomenclature.